Let f(x) = x/(x - 1), g(x) = (x - 1)/x, h(x) = 1/(x - 1) .
Then (gh)(x) =
a. 1
b. x
c. 1/x
d. -x
e. none of these

Answer by stanbon(75887)

Let f(x) = x/(x - 1), g(x) = (x - 1)/x, h(x) = 1/(x - 1) .
Then (gh)(x) =g(1/(x-1)
=[(1/x-1)-1)/(1/x -1)]
=[(1-(x-1)/(x-1)]/[1/(x-1)]
=[(2-x)/(x-1)]
=2-x
Answer; None of these
Cheers,
Stan H.